OpenStudy (anonymous):

Evaluate if possible the function (if possible) at the given values of the independant variables F(x)=X^3-X F(X)-F(1)/X-1

OpenStudy (anonymous):

X^3-X - (1^3 -1) /(X-1) X^3 -X -0 /(X-1) X(X^2 - 1)/(X-1) X(X+1) (X-1) /(X-1) X(X+1) = X^2 + X

OpenStudy (anonymous):

F(X)=X^3-X F(1)=1-1=0 so: (F(X)-F(1))/X-1= (X^3-X-0)/(X-1) simplifies to ^ that answer

OpenStudy (anonymous):

remember this problem when you get to calculus and learn about the difference quotient. This new function tells you the slope of the secant line from X to 1 of the graph of X^3-X for x not equal 1. It will be very useful when you get further along. :)